# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ 4 On the completeness of invariant geometric models
In this section, we move on to demonstrate and prove the E(3)-completeness of a broad range of invariant models that leverage more geometric features or more complicated aggregation schemes than DisGNN. These models can effectively identify all those highly-symmetric point clouds that DisGNN can not. We first introduce a new simple invariant model design, which can precisely break such symmetry and achieve completeness. Consequently, we show how this new design serves as a theoretical tool for proving a wide range of invariant models’ expressiveness.
@@ 4.1 GeoNGNN: breaking symmetry through node marking
According to Theorem 3.2 (along with the lemmas emphasized in its key idea), the essential factor in identifying a point cloud by DisGNN is the existence of two distinct anchors. Although every point cloud inherently possesses one, i.e., the geometric center, the other may not exist in some highly symmetric cases, such as a sphere. This hinders DisGNN from achieving completeness. In this subsection, we show that GeoNGNN, the geometric counterpart of Nested GNN (NGNN) [63], can fill in this last piece of the puzzle by applying DisGNN on point clouds with an additionally marked node, which breaks the symmetry and exactly acts as the other anchor.
NGNN [63] is a subgraph GNN first proposed in traditional graph learning. It performs a two-level GNN on a given graph, where the first level, known as the inner GNN, runs concurrently on each node’s $k$-hop ego subgraph with the center node optionally marked (gives it a distinct feature), summarizing the local subgraph information as the nodes’ initial embedding in the original graph, and the second level, referred to as the outer GNN, runs on the original graph to obtain the final node and graph representations. We now extend NGNN to geometric settings as follows:
Definition 4.1.
(GeoNGNN) GeoNGNN is the geometric counterpart of NGNN [63], which takes point clouds $P\in\mathbb{R}^{n\times 3}$ (potentially with node features $X\in\mathbb{R}^{n\times d}$, such as atomic numbers) as input, treats it as a distance graph with cutoff $r_{\text{cutoff}}$, and
  • Both the inner and outer GNN adopt DisGNN, with $N_{\text{in}}$ and $N_{\text{out}}$ layers each.
  • The ego subgraph of each node $i$ contains all nodes and edges within Euclidean distance cutoff $r_{\text{sub}}$, with node $i$ being explicitly marked.
Overall, GeoNGNN adopts the same framework as NGNN, except that it is tailored for geometric settings: the subgraph cutoff is defined on Euclidean distance instead of “hop”, while the base model is the simplest invariant model, DisGNN. We are now ready to prove the completeness of GeoNGNN:
Theorem 4.2.
(E(3)-Completeness of GeoNGNN) When the following conditions are met, GeoNGNN is E(3)-complete:
@@
Key idea
Given an arbitrary point cloud $P$, consider a specific node $i$ within it that differs from the geometric center. As node $i$ is explicitly marked within its own subgraph, there are now two anchors in this subgraph, namely node $i$ and the geometric center, that facilitate triangular distance encoding by DisGNN. Consequently, DisGNN can fully represent and identify $P$ through the representation of node $i$’s subgraph, as underscored in the key idea of Theorem 3.2. Notice that in a point cloud with more than two nodes, a node that is distinct from the geometric center always exists, and through the outer pooling, overall completeness and permutation invariance are guaranteed.
While Theorem 4.2 establishes that infinite subgraph radius and distance cutoff guarantee completeness over all point clouds, in Appendix C.2 we show that finite values also lead to boosted expressiveness compared to DisGNN. And we notice that the conditions specified in Theorem 4.2 ultimately result in polynomial-time complexity w.r.t. the size of the point cloud even when considering the complexity involved in obtaining intermediate injective functions, which aligns with prior studies [35, 59, 26, 37, 10]. The detailed complexity analysis is provided in Appendix C.1.
@@ 4.2 Completeness of geometric subgraph GNNs
GeoNGNN provides a simple approach to extend traditional subgraph GNNs to geometric scenarios with remarkable geometric expressiveness. Indeed, the realm of traditional graph learning literature contains a multitude of subgraph GNNs such as DSS-GNN [5], GNN-AK [64], OSAN [49] and so on [61, 17]. Extending these models to geometric settings can significantly enlarge the design space of geometric models, and potentially introduce valuable inductive biases. In this section, we take a pioneering step to do this, subsequently establishing the geometric completeness of all these models.
We first formally define general geometric subgraph GNN, which represents a large family of geometric models. The definition extends the general subgraph GNNs defined in the seminal work [62], which operate on unweighted graphs, to geometric scenarios wherein they operate on point clouds. We briefly introduce modifications here, while self-contained definitions can be found in Appendix D.
Definition 4.3.
(General geometric subgraph GNN, informal) A general geometric subgraph GNN takes point clouds $P\in\mathbb{R}^{n\times 3}$ (potentially with node features $X\in\mathbb{R}^{n\times d}$, such as atomic numbers) as input, treats it as a distance graph with cutoff $r_{\text{cutoff}}$, and:
  • It employs node marking with $r_{\text{sub}}$-size ego subgraph as the subgraph generation policy.
  • It stacks multiple geometric subgraph GNN layers, each of which has the form of Definition 2.1 in [62], with the only modifications on local operations: The two local operations for updating $h_{uv}$ (node $v$’s representation in $u$’s subgraph) additionally integrate distance information, which are represented by $\{\!\!\{\big{(}h_{uw},d_{vw}\big{)}\mid w\in\mathcal{N}(v)\}\!\!\}$ and $\{\!\!\{\big{(}h_{wv},d_{uw}\big{)}\mid w\in\mathcal{N}(u)\}\!\!\}$.
  • It adopts vertex-subgraph or subgraph-vertex pooling schemes [62] to obtain graph representation.
@@
Theorem 4.4.
(Completeness for general geometric subgraph GNNs) Under the conditions specified in Theorem 4.2, all general geometric subgraph GNNs in Definition 4.3 with at least one local aggregation are E(3)-complete.
Key idea
It can be shown that any type of general geometric subgraph GNNs, which could employ any combination of aggregation and pooling schemes, can implement [17] GeoNGNN, which is established as an E(3)-complete model in Theorem 4.2. As a consequence, any general geometric subgraph GNN can also achieve E(3)-completeness under the same conditions as GeoNGNN.
Concerning prior research on traditional subgraph GNNs [5, 64, 49, 61, 17], their geometric counterparts can be defined correspondingly by substituting all local aggregation schemes with distance-aware counterparts. We denote their geometric counterpart by prefixing Geo to their original names, for instance, GeoSUN [17]. Consequently, it can be established that some of them precisely fall into the general definition 4.3, such as GeoOSAN [49], and for the others such as GeoGNN-AK [64] and GeoSUN [17], while even though they do not exactly match the general definition, they can still implement GeoNGNN, thereby exhibiting E(3)-completeness (Theorem D.2). Please refer to Appendix D.2 for details about all of these.
We note that all subgraph GNNs discussed in this section have been proven to be strictly weaker than 2FWL [7] in traditional graph learning context [62]. Specifically, there exist pairs of unweighted graphs that can be distinguished by 2FWL but remain indistinguishable by these subgraph GNNs. Moreover, these subgraph GNNs themselves establish a strict expressiveness hierarchy [62]. Interestingly, our findings reveal that all of these discrepancies diminish when these models are extended to geometric scenarios with point clouds by leveraging distance graphs. This phenomenon could be attributed to the low-rank nature of distance graphs, whose intrinsic dimension is less than the point cloud space dimension $3n$ rather than equals to their ambient dimension $n^{2}$.
@@ 4.3 Completeness of well-established invariant models
Based on the completeness of GeoNGNN, we move on to establish the E(3)-completeness of several well-established invariant models, including DimeNet [20], GemNet [22] and SphereNet [39]. These models do not exactly learn on subgraphs, however, can still be mathematically aligned with GeoNGNN. Descriptions of these models and proofs can be found in Appendix E.5.
Theorem 4.5.
(E(3)-Completeness of DimeNet, SphereNet, GemNet) When the following conditions are met, DimeNet, SphereNet [^3] Here in SphereNet we do not consider the relative azimuthal angle $\varphi$, since SphereNet with $\varphi$ is not E(3)-invariant, while E(3)-completeness is defined on E(3)-invariant models. Note that the exclusion of $\varphi$ only results in weaker expressiveness. and GemNet are E(3)-complete.
  • The aggregation layer number is larger than some constant $C$, which is irrelevant to the node number.
@@
Key idea
The key insight underlying is that all these models track edge representations ($h_{ij}^{\text{edge}}$ for edge $(i,j)$), which can be mathematically aligned with the node-subgraph representations tracked in GeoNGNN ($h_{ij}^{\text{subg}}$ for node $i$ in subgraph $j$). Moreover, the additionally incorporated features, such as angles, can all be equivalently expressed by multiple distances. Consequently, these three models can all implement GeoNGNN, and thereby achieving completeness.
Notice that DimeNet is a relatively simpler invariant model that aggregates neighbor information in a weaker manner compared to existing 2FWL-like complete geometric models such as 2F-DisGNN [37]. Specifically, when updating $h_{ij}$, DimeNet can be equivalently considered as aggregating $(h_{ki},d_{kj})$ for a specific neighbor $k$, while 2F-DisGNN aggregates $(h_{ik},h_{kj})$. The latter incorporates a more informative hidden representation $h_{kj}$, which is essential in the proof of 2F-DisGNN’s completeness [37, 10], instead of $d_{kj}$ in the former. Nevertheless, our findings for the first time show that these models are equally E(3)-complete under similar conditions (i.e., fully connected distance graphs). Consequently, future developments like GemNet, which integrate higher-order geometry, could be unnecessary in terms of boosting theoretical expressiveness under such conditions.
In summary of this entire section, we have characterized a broad collection of E(3)-complete invariant models under the similar conditions. However, it is essential to acknowledge that while two models may exhibit equivalence in terms of completeness, they may encompass distinct inductive biases that could lead to differing levels of generalization capability and experimental performance, as exemplified in Li et al. [37] and in our experiment sections.
@@ 5 Experiments
In this section, we additionally evaluate GeoNGNN, the simplest complete model within the large family we established, to assess its practical performance. We specifically focus on its capacity to effectively distinguish symmetric point clouds in practical scenarios where numerical precision could potentially impact the outcomes despite the theoretical assurance of completeness. Furthermore, we also explore the performance of GeoNGNN in molecule-relevant downstream tasks, where the capability for geometric learning is essential for accurately approximating target functions.
We compare GeoNGNN with two main types of geometric models. The first type integrates high-order SO(3) irreducible representations (including 1-order, i.e., vectors) and mostly employs tensor products to produce equivariant outputs. These models are referred to as equivariant models, including PaiNN [52], NequIP [4] and MACE [3]. The second type integrates high-order invariant geometry, such as angle, dihedral angle, or learned representations. We refer to these models as invariant models, including DimeNet [20], GemNet [22] and 2-F-DisGNN [37]. Note that all of these models can capture high-order geometry, enabling them to achieve state-of-the-art performance on downstream tasks. Given that GeoNGNN can be regarded as the subgraph-enhanced version of DisGNN, we also implement and evaluate DisGNN, which includes more complicated residual layers than its special case SchNet [53], as a baseline to assess the direct improvement achieved through subgraph representation. Detailed experimental settings are provided in Appendix G.
[…]
Equivariant models Invariant models
Molecule Target PaiNN NequIP MACE DimeNet GemNet 2F-Dis. Vanilla DisGNN GeoNGNN
[…]
Naphthalene F 0.0830 $0$ul0.0369 0.1105 0.0438 0.0478 0.4144 0.0377
[…]
AVG RANK 6.71 2.80 2.50 5.70 5.00 2.25 6.55 2.55
5.1 Separation power: synthetic counterexamples

### v3

@@ 5 On the Completeness of Invariant Geometric Models
In this section, we move on to demonstrate and prove the E(3)-completeness of a broad range of invariant models that leverage more geometric features or more complicated aggregation schemes than DisGNN under fully-connected conditions. These models can effectively identify all those highly-symmetric point clouds that DisGNN can not. We first introduce a new simple invariant model design, which can precisely break such symmetry and achieve completeness. Consequently, we show how this new design serves as a theoretical tool for proving a wide range of invariant models’ expressiveness. Proof for this section is provided in Appendix B.
@@ 5.1 GeoNGNN: Breaking Symmetry Through Node Marking
According to Theorem 4.2 (along with the lemmas emphasized in its key idea), the essential factor in identifying a point cloud by DisGNN is the existence of two distinct anchors. Although every point cloud inherently possesses one, i.e., the geometric center, the other may not exist in some highly symmetric cases, such as a sphere. This hinders DisGNN from achieving completeness. In this subsection, we show that GeoNGNN, the geometric counterpart of Nested GNN (NGNN) (Zhang & Li, 2021) , can fill in this last piece of the puzzle by applying DisGNN on point clouds with an additionally marked node, which breaks the symmetry and exactly acts as the other anchor.
GeoNGNN employs a two-level hierarchical framework by nesting DisGNN. Specifically, it processes a given point cloud as follows: 1) The first level, referred to as the inner GNN, operates independently on each point’s $r_{\text{sub}}$-sized sub-point cloud and aggregates the local sub-point cloud information into an initial embedding for the corresponding point in the original point cloud. 2) The second level, referred to as the outer GNN, processes the original point cloud based on the embeddings generated by the inner GNN, thereby producing final point-level and cloud-level representations.
Formally, for the inner GNN, the representation of the node $j$ in node $i$’s sub-point cloud at the $l$-th layer, denoted as $h_{ij}^{(l)}$, is updated as:
$h_{ij}^{(l+1)}=f_{\text{update, inner}}^{(l)}\left(h_{ij}^{(l)},f_{\text{aggr,inner}}^{(l)}\left(\{\!\!\{(h_{ik}^{(l)},d_{kj})\mid k\in\mathcal{N}(j),d_{ik}\leq r_{\text{sub}}\}\!\!\}\right)\right),$ (2)
for all $h_{ij}$ satisfying $d_{ij}\leq r_{\text{sub}}$, and $\mathcal{N}(j)$ represents all nodes $k$ satisfying $d_{kj}\leq r_{\text{cutoff}}$. Here, $r_{\text{sub}}$ and $r_{\text{cutoff}}$ are hyperparameters representing the subgraph size and interaction cutoff, respectively. $h_{ij}^{(0)}$ is initialized as:
$h_{ij}^{(0)}=f_{\text{init, inner}}(x_{j},d_{ij},\mathds{1}_{i=j}),$ (3)
where $x_{j}$ is the raw feature of node $j$, and $d_{ij}$ and $\mathds{1}_{i=j}$ denote the position encoding and marking of the center point, respectively. After $N_{\text{in}}$ iterations, the inner GNN summarizes the sub-point cloud as:
$h_{i}^{(0)}=f_{\text{output, inner}}(\{\!\!\{h_{ij}^{(N_{\text{in}})}\mid j\in[n],d_{ij}\leq r_{\text{sub}}\}\!\!\}),$ (4)
producing the initial point representations for the outer GNN. The outer GNN then updates these representations over $N_{\text{out}}$ iterations following the framework in Equation 1.
By representing point clouds as distance graphs, GeoNGNN captures subgraph patterns instead of subtree patterns, which have been shown to be more expressive in traditional graph learning (Zhang & Li, 2021). We now show that GeoNGNN achieves geometric completeness under fully-connected conditions, thereby effectively addressing expressiveness limitations of DisGNN.
Theorem 5.1 (E(3)-Completeness of GeoNGNN).
When the following conditions are met, GeoNGNN is E(3)-complete:
@@
Key idea. Given an arbitrary point cloud $P$, consider a specific node $i$ within it that differs from the geometric center. As node $i$ is explicitly marked within its own subgraph (Equation 3), there are now two anchors in this subgraph, namely node $i$ and the geometric center, that facilitate triangular distance encoding by DisGNN. Consequently, DisGNN can fully represent and identify $P$ through the representation of node $i$’s subgraph, as underscored in the key idea of Theorem 4.2. Notice that in a point cloud with more than two nodes, a node that is distinct from the geometric center always exists, and through the outer pooling, overall completeness and permutation invariance are guaranteed.
While Theorem 5.1 establishes that infinite subgraph radius and distance cutoff guarantee completeness over all point clouds, in Appendix D.2 we show that finite values also lead to boosted expressiveness compared to DisGNN. And we notice that the conditions specified in Theorem 5.1 ultimately result in polynomial-time complexity w.r.t. the size of the point cloud even when considering the complexity involved in obtaining intermediate injective functions, which aligns with prior studies (Kurlin, 2023; Widdowson & Kurlin, 2023; Hordan et al., 2024b; Li et al., 2024; Delle Rose et al., 2023). The detailed complexity analysis is provided in Appendix D.1.
@@ 5.2 Completeness of Geometric Subgraph GNNs
GeoNGNN provides a simple approach to extend traditional subgraph GNNs to geometric scenarios with remarkable geometric expressiveness. Indeed, the realm of traditional graph learning literature contains a multitude of subgraph GNNs such as DSS-GNN (Bevilacqua et al., 2021), GNN-AK (Zhao et al., 2021), OSAN (Qian et al., 2022) and so on (You et al., 2021; Frasca et al., 2022). Extending these models to geometric settings can significantly enlarge the design space of geometric models, and potentially introduce valuable inductive biases. In this section, we take a pioneering step to do this, subsequently establishing the geometric completeness of all these models.
We first define the general geometric subgraph GNN, a broad family of geometric models, by slightly adapting the unweighted graph models from Zhang et al. (2023) to handle geometric scenarios. For simplicity, we provide an intuitive overview and outline the modifications here, while self-contained and formal definitions can be found in Appendix E.
Definition 5.2 (General geometric subgraph GNN, informal).
A general geometric subgraph GNN takes point clouds $P\in\mathbb{R}^{n\times 3}$ (potentially with node features $X\in\mathbb{R}^{n\times d}$, such as atomic numbers) as input, treats it as a distance graph with interaction cutoff $r_{\text{cutoff}}$, and:
  • Utilizes node marking with $r_{\text{sub}}$-size ego subgraph as the subgraph generation policy.
  • Stacks multiple geometric subgraph GNN layers, each following general subgraph GNN layer defined in Zhang et al. (2023), which could arbitrarily include single-point, global, and local operations for aggregating graph information at different levels. The only modification is to the local operations: when updating $h_{uv}$ (the representation of node $v$ in $u$’s subgraph), distance information is additionally integrated, thus the local operations aggregate information $\{\!\!\{\big{(}h_{uw},d_{vw}\big{)}\mid w\in\mathcal{N}(v)\}\!\!\}$ and $\{\!\!\{\big{(}h_{wv},d_{uw}\big{)}\mid w\in\mathcal{N}(u)\}\!\!\}$ respectively.
  • It adopts vertex-subgraph or subgraph-vertex pooling schemes (Zhang et al., 2023) to summarize node and subgraph features and produce the final graph representation.
@@
Theorem 5.3 (Completeness for general geometric subgraph GNNs).
Under the conditions specified in Theorem 5.1, all general geometric subgraph GNNs in Definition 5.2 with at least one local aggregation are E(3)-complete.
Key idea. It can be shown that any type of general geometric subgraph GNNs, which could employ any combination of aggregation and pooling schemes, can implement (Frasca et al., 2022) GeoNGNN, which is established as an E(3)-complete model in Theorem 5.1. Consequently, any general geometric subgraph GNN can also achieve E(3)-completeness under the same conditions as GeoNGNN.
Concerning prior research on traditional subgraph GNNs (Bevilacqua et al., 2021; Zhao et al., 2021; Qian et al., 2022; You et al., 2021; Frasca et al., 2022), their geometric counterparts can be defined correspondingly, by substituting all local aggregation schemes with distance-aware counterparts. We denote their geometric counterpart by prefixing Geo to their original names, for instance, GeoSUN Frasca et al. (2022). Consequently, it can be established that some of them precisely fall into the general definition 5.2, such as GeoOSAN (Qian et al., 2022), and for the others such as GeoGNN-AK (Zhao et al., 2021) and GeoSUN Frasca et al. (2022), while even though they do not exactly match the general definition, they can still implement GeoNGNN, thereby exhibiting E(3)-completeness (Theorem E.2). Please refer to Appendix E.2 for details about all of these.
We note that all subgraph GNNs in this section have been proven to be strictly weaker than 2-FWL (Cai et al., 1992) in traditional graph learning context (Zhang et al., 2023). Specifically, there exist pairs of unweighted graphs that can be distinguished by 2-FWL but remain indistinguishable by these subgraph GNNs. Moreover, these subgraph GNNs themselves establish a strict expressiveness hierarchy (Zhang et al., 2023). Interestingly, our findings reveal that all of these discrepancies diminish when these models are extended to geometric scenarios with point clouds by leveraging distance graphs. This phenomenon could be attributed to the low-rank nature of distance graphs, whose intrinsic dimension is less than the point cloud space dimension $3n$ rather than equals to their ambient dimension $n^{2}$.
@@ 5.3 Completeness of Well-Established Invariant Models
Based on the completeness of GeoNGNN, we move on to establish the E(3)-completeness of several well-established invariant models, including DimeNet (Gasteiger et al., 2019), GemNet (Gasteiger et al., 2021) and SphereNet (Liu et al., 2021). These models do not exactly learn on subgraphs, however, can still be mathematically aligned with GeoNGNN. Formal descriptions of these models can be found in Appendix B.6.
Theorem 5.4 (E(3)-Completeness of DimeNet, SphereNet, GemNet).
When the following conditions are met, DimeNet, SphereNet [^4] Here in SphereNet we do not consider the relative azimuthal angle $\varphi$, since SphereNet with $\varphi$ is not E(3)-invariant, while E(3)-completeness is defined on E(3)-invariant models. Note that the exclusion of $\varphi$ only results in weaker expressiveness. and GemNet are E(3)-complete.
  • The aggregation layer number is larger than some constant $C$ (irrelevant to the node number).
@@
Key idea. The key insight underlying is that all these models track edge representations ($h_{ij}^{\text{edge}}$ for edge $(i,j)$), which can be mathematically aligned with the node-subgraph representations tracked in GeoNGNN ($h_{ij}^{\text{subg}}$ for node $j$ in subgraph $i$). Moreover, the additionally incorporated features, such as angles, can all be equivalently expressed by multiple distances. Consequently, these three models can all implement GeoNGNN, and thereby achieving completeness.
Note that DimeNet is a relatively simpler invariant model that aggregates neighbor information in a weaker manner compared to existing 2-FWL-like complete geometric models such as 2F-DisGNN (Li et al., 2024). Specifically, when updating $h_{ij}$, DimeNet can be equivalently considered as aggregating $(h_{ki},d_{kj})$ for a specific neighbor $k$, while 2F-DisGNN aggregates $(h_{ik},h_{kj})$. The latter incorporates a more informative hidden representation $h_{kj}$, which is essential in the proof of 2F-DisGNN’s completeness (Li et al., 2024; Delle Rose et al., 2023), instead of $d_{kj}$ in the former. Nevertheless, our findings first show that these models are equally E(3)-complete under fully-connected conditions. Consequently, future developments like GemNet that integrate higher-order geometry could be unnecessary in terms of boosting theoretical expressiveness under such conditions.
5.4 Summarization and Discussion
Fully-Connected Condition. Thus far, we have characterized a broad collection of E(3)-complete invariant models under similar conditions. Among these conditions, the condition of fully-connected geometric graphs (or equivalently, infinite cutoffs) is required. Notably, this condition is consistent with all prior works that rigorously characterize invariant models/descriptors in the same sense, including (Kurlin, 2023; Widdowson & Kurlin, 2022; 2023; Li et al., 2024; Delle Rose et al., 2023; Hordan et al., 2024b; a). And due to significant local information loss during invariant message passing (Joshi et al., 2023; Du et al., 2024), even under this condition the characterization remains highly nontrivial. Indeed, removing the fully-connected condition would require conditions about specific forms of sparsity, which would be considerably more demanding than the overall connectivity typically required by equivariant models (Joshi et al., 2023; Du et al., 2024; Wang et al., 2024; Sverdlov & Dym, 2024) that can maintain local information through equivariant features. Such characterization is left for future work.
Theoretical Characterization vs. Practical Use. In practical scenarios, efficiency and generalization are often prioritized, and local/sparse connectivity can typically offer empirical advantages in these aspects (Musaelian et al., 2023). Thus, strictly adhering to the complete condition which requires full connectivity is unnecessary, and finding a balance is crucial, as demonstrated in our additional experiments in Appendix F (Figure 6). However, this does not diminish the importance of theoretical characterization, since theoretical completeness provides an upper bound for the parametric model’s potential–when combined with MLPs (Hornik et al., 1989), complete models can achieve universal approximation over continuous invariant functions (Hordan et al., 2024b; Li et al., 2024). This is analogous to the universal approximation property for MLPs (Hornik et al., 1989) and the Turing completeness for RNNs (Siegelmann & Sontag, 1992), where practical implementations do not (and typically cannot) satisfy all theoretical conditions, yet these results indicate their superiority over weaker structures that cannot achieve certain approximations even with unlimited resources.
SE(3)-complete Counterpart. Finally, we present a provable SE(3)-complete variant of GeoNGNN, capable of distinguishing chiral molecules, by making a minor modification to the distance features through the addition of a orientation sign. This is further detailed in Appendix H.
@@ 6 Experiments
In this section, we conduct additional assessments to validate our theoretical claims. The first experiment aims to verify the conclusion that DisGNN is nearly complete by assessing the proportion of its unidentifiable cases in real-world point clouds. The second experiment is designed to evaluate whether the complete models consistently demonstrate separation power for challenging pairs of point clouds, where numerical precision may influence the outcomes. We provide more evaluations of GeoNGNN on practical molecular-relevant tasks in Appendix F, which could offer further insights.
6.1 Assessment of Unidentifiable Cases of DisGNN
In Theorem 4.3, we have rigorously shown the rarity of symmetric point cloud sets, which are supersets of the unidentifiable set of DisGNN. However, since real-world point clouds are typically subject to slight noise, requiring exact symmetry would be overly restrictive. Therefore, we address a more challenging setting by explicitly accounting for noise and evaluating the rarity of relaxed symmetric point clouds in practical scenarios. To this end, we first define two noise tolerances that allow us to relax the exact $\mathcal{C}$- and $\mathcal{D}$-symmetry.
Noise Tolerances. 1) The rounding number $r$ for distance-related calculations. When applying algorithms $\mathcal{C}$ and $\mathcal{D}$, we round the distance values to $r$ decimal places for robustness to noise. 2) The deviation error $\epsilon$. In Definition 4.1, we define symmetry by the set of extended "mass" centers. We provide an equivalent definition in Proposition C.4, which shows that a point cloud is $\mathcal{A}$-symmetric if and only if the geometric centers of all sub-point clouds, partitioned by distinct node features of $\mathcal{A}$, coincide. This alternative definition allows us to define another noise tolerance: we now consider a point cloud as symmetric if all sub-point clouds’ centers described above lie within a ball of radius $\epsilon$. Together, $r$ and $\epsilon$ define a relaxed symmetry that accounts for noise in real-world point clouds. When $r\to+\infty$ and $\epsilon\to 0$, the relaxed symmetry becomes the exact theoretical symmetry.
We are now ready to evaluate the proportion of relaxed symmetric point clouds in real-world datasets. We select two representative datasets, namely QM9 (Ramakrishnan et al., 2014; Wu et al., 2018) and ModelNet40 (Wu et al., 2015), for this assessment. The QM9 dataset comprises approximately 130K molecules, while the ModelNet40 dataset consists of roughly 12K real-world point clouds categorized into 40 classes, including objects such as chairs. We first rescale all point clouds, and then fix the rounding number to small values and evaluate the proportion of symmetric point clouds with respect to different values of $\epsilon$. Please refer to Section C.3 for detailed settings.
Results are shown in Figure 2. Here are several key observations: 1) For the QM9 dataset, even with the largest deviation error ($10^{-1}$), the proportion of ${\mathcal{C}}$- and ${\mathcal{D}}$-symmetric point clouds is less than $\sim$0.15of the entire dataset. Since ${\mathcal{C}}$- and ${\mathcal{D}}$-symmetric point cloud sets are supersets of the unidentifiable set of DisGNN, the unidentifiable proportion is therefore no more than $\sim$0.15%. With the strictest deviation error, only 0.0046% of the graphs (6 out of $\sim$130K) exhibit ${\mathcal{D}}$-symmetry. 2) For the ModelNet40 dataset, when the deviation error is less than $10^{-2}$, only 1 point cloud exhibit $\mathcal{C}$-symmetry. Moreover, no ${\mathcal{D}}$-symmetric point clouds are found across all deviation errors.
To summarize, the statistical results suggest that the occurrence of unidentifiable graphs in DisGNN is practically negligible in real-world scenarios, even when the criterions are significantly relaxed. This supports the conclusion that DisGNN is almost E(3)-complete.
Figure 2: Assessment of symmetric point clouds in real-world datasets. (a) Proportion of symmetric point clouds in QM9 ($r=2$). (b) Proportion of symmetric point clouds in ModelNet40. The proportion of ${\mathcal{D}}$-symmetric point clouds in ModelNet40 is zero across all deviation errors ($r=1$).
6.2 Separation Power on Synthetic Point Cloud Pairs